The shelf life of a particular dairy product is normally distributed with a mean of 12 days and a standard deviation of 3 days. About what percent of this product last between 10 days and 13 days? Do not use the empirical rule. Use the graphing calculator, not a table, to find the answer to one decimal place.
Answer #1

Solution:-

P(10 < X < 13) = P((10-12)/3 < (X-mean)/sd < (13-12)/3 )
= P(-0.6666 < Z < 0.3333)
= P(Z < 0.3333) < P(Z < -0.6666)
= 0.6293 - 0.2514
= 0.3779

The probability that -0.6666<2<0.3333 is equal to the blue area under the curve. 0.30 У 0.30 0.10 0.08 0.00 -0.67 0.33 -0.03